the sum of two number is 8, and the sum of their squares is 34. what is the smaller number
Answer by Edwin McCravy(20081)

the sum of two number is 8, and the sum of their squares is 34. what is the smaller number
                x + y  =  8
               x² + y² = 34

Solve the first equation for y

                     y = 8 - x

Substitute (8 - x) for y in the second equation:

         x² + (8 - x)² = 34 
   x² + (8 - x)(8 - x) = 34
x² + 64 - 8x - 8x + x² = 34
        2x² + 64 - 16x = 34

Get 0 on the right:

        2x² + 30 - 16x = 0

Attange in order of descending powers 

        2x² - 16x + 30 = 0

Divide every term by 2

          x² - 8x + 15 = 0

Factor the left side:

        (x - 5)(x - 3) = 0

Use the zero-factor principle:

x - 5 = 0;     x - 3 = 0
    x = 5;         x = 3

We must find the corresponding value of y for
each of these solutions for x:

for x = 5, substitute in

                     y = 8 - x
                     y = 8 - 5
                     y = 3

So one solution is (x,y) = (5,3)

for x = 3, substitute in

                     y = 8 - x
                     y = 8 - 3
                     y = 5

So the other solution is (x,y) = (3,5)

The question asked for the smaller number, which is 3.

FYI, the graphical representation of the system is below

The equation x² + y² = 34 has a graph which is a circle, 
and x + y = 8 has a graph which is a straight line and 
the line cuts the circle in two points (3,5) and (5,3), 
as you see below:


Edwin